# IP-CRR: Information Pursuit for Interpretable Classification of Chest Radiology Reports

### v1

@@ 4.1 Dataset and Implementation Details
Dataset. To verify the effectiveness of the proposed method, we conduct experiments on a public radiology dataset named MIMIC-CXR [15]. In this dataset, each patient is called “subject”, and each examination is called “study”. Each study corresponds to one radiology report and multi-label ground-truth label indicating whether diseases are present. The dataset is randomly split into 132296, 14914, and 39613 studies separately for training, validation, and testing.
Implementation Details. To train the interpretable model, a querier and a classifier are designed. The architecture of these two components are both five-layer multi-layer perceptions (see code for details). The IP-CRR model terminates with a confidence threshold of 0.85 or a maximum number of queries of 200. The optimizer is AdamW [20] with an initial learning rate of 0.0001 and a cosine annealing decay. The model is trained with batch size 128 and 1000 epochs. The loss function is weighted binary cross-entropy to alleviate the class imbalance. All experiments are implemented in PyTorch 2.5.1 and on an NVIDIA A5000 GPU with a memory of 24.5GB.
@@
(a) Average precision over number of queries asked. Dashed line represents the maximum performance over the number of queries.
@@
Figure 3: Quantitative evaluation of IP-CRR as the number of queries increase.
@@ 4.2 Quantitative Results and Comparisons
We evaluate the predictive performance of IP-CRR with all queries on six binary prediction tasks: Lung Opacity (LO), Calcification of the Aorta (CA), Support Devices (SD), Cardiomegaly (CM), Pleural Effusion (PE), and Pneumonia (PN). Our baseline methods include 1) CXR-BERT (FT-Last): CXR-BERT [3], with fine-tuning on the last layer; 2) CXR-BERT (FT-All): CXR-BERT [3], with fine-tuning on all layers; 3) Flan-T5-large model [11], without any fine-tuning; and 4) CBM [18]. Notably, 1-3) are in the class of black-box models and 4) is an interpretable-by-design model. The results are listed in Figure 1. Our metric of choice is the Average Precision (AP) and F1 score.
Compared with black-box models, IP-CRR achieves a higher AP and F1 score against CXR-BERT (FT-Last), Flan-T5-large, and CBM for LO, CA, CM, and PN, and achieves comparable performance for SD. However, when compared with CXR-BERT (FT-All), IP-CRR achieves a similar AP for LO, but falls short on other tasks. We believe the difference in performance is largely attributed to CXR-BERT’s black-box nature, its complex architecture, and the fact that it was fine-tuned on BERT. On the contrary, IP-CRR is implemented using a simple MLP with much fewer parameters.
Furthermore, in Figure 3, we evaluate the overall AP as a function of the number of queries. For the task of LO, IP-CRR reaches nearly 0.95 AP with just 30 queries (out of 520 total), and improves to 0.97 with 100 queries – less than 20% of the total number of queries. We find a similar trend for other tasks, which demonstrates that IP-CRR generally selects informative queries in its early iterations. Alternatively, we observe that IP-CRR converges to the maximum performance it can achieve with only a small subset of queries. Compared to baselines in Table 1, IP-CRR achieves a better performance with smaller number of queries than CBMs with all available ones.
Last but not least, in Figure 3, we plot the number of queries needed as the threshold on the posterior changes. The threshold on the posterior approximately measures the confidence the user requires the predictor to reach before terminating the IP-CRR algorithm. We observe that the IP-CRR predictor requires fewer than 150 queries to achieve a confidence of 0.95 for most of tasks. This demonstrates that IP-CRR is able to find informative queries to make confident predictions.
@@ 4.3 Qualitative Analysis of Query-answer Chains
Figure 4: Examples of query-answer chains from IP-CRR. Each row in the matrix represents the posterior $P(Y\mid q_{1:k}(x))$ after selecting the query $q_{k}$. The text is $q_{k}$, with red implying the fact is not present $q_{k}(x)=-1$, green implying is present $q_{k}(x)=1$, and black implying it cannot be answered $q_{k}(x)=0$. “N” or “P” means negative or positive prediction of the disease, respectively.
The core mechanics of using IP-CRR for interpretable predictions is the iterative process of selecting informative queries. This allows the user to observe how the posterior evolves as the number of queries increases. In Figure 4, we demonstrate two examples of query-answer chains obtained using IP-CRR. For the example on the left, our method first selects $q_{1}=$Opacification likely reflects atelectasis?”. From the report, the model infers that $q_{1}(x)$ is not present in the report. Then, it selects the query $q_{2}=$Patchy opacity at right lung base?”. Continuing in this fashion, after 9 queries, the model concludes that the patient has LO. The same sequential deduction can be observed in the example on the right, highlighting the interpretability IP-CRR offers when applied to radiology reports.
@@ 5 Conclusion
In this work, we propose an interpretable-by-design text classification framework. We obtain a large set of queries by mining existing radiology reports, and leverage NLI for answering these queries. To evaluate the proposed framework, we conducted experiments on a large-scale radiology dataset and on multiple tasks, with both quantitative and qualitative results that illustrate interpretability and competitive performance. As for limitation, our current method relies on ternary query-answers for all queries, and we reserve leveraging stronger Foundation Models in the medical domain for more expressive query-answers as future work.
Acknowledgments
Yuyan, Ryan, and René acknowledge the support from the Penn Engineering Dean’s Fellowship, the startup funds, and the NIH grant R01NS135613. Pablo received support in Chile from the National Agency for Research and Development (ANID) through the Millennium Institute for Intelligent Healthcare Engineering - iHEALTH (ICN2021_004), the National Center for Artificial Intelligence - CENIA (FB210017), and the ANID Scholarship Program / Doctorado Becas Chile / 2019 - 21191569.

### v2

@@ 4.1 Dataset and Implementation Details
Dataset. We evaluate the proposed method on a public radiology dataset named MIMIC-CXR [15]. In this dataset, each patient is called “subject”, and each examination is called “study”. Each study has one radiology report and multi-label ground-truth annotations [13] indicating whether diseases are present. The dataset is randomly split into training, validation, and test sets in a 7:1:2 ratio.
Implementation Details. The architecture of both the querier and the classifier consists of a multilayer perceptron with 5 layers (see code for details). The IP-CRR model terminates with a confidence threshold of 0.85 or a maximum number of queries of 200. The optimizer is AdamW [20] with an initial learning rate of 0.0001 and a cosine annealing decay. The model is trained with a batch size of 128 for 1000 epochs. The loss function is weighted binary cross-entropy to alleviate the class imbalance. All experiments are implemented in PyTorch 2.5.1 using an NVIDIA A5000 GPU with a memory of 24.5GB.
@@
(a) Average precision versus number of queries asked. Dashed lines mark the maximum performance over all queries.
@@
Figure 3: Quantitative evaluation of IP-CRR as the number of queries increases.
@@ 4.2 Quantitative Results and Comparisons
We evaluate the performance of IP-CRR with all queries on six binary classification tasks: Lung Opacity (LO), Calcification of the Aorta (CA), Support Devices (SD), Cardiomegaly (CM), Pleural Effusion (PE), and Pneumonia (PN). Baselines include CXR-BERT [3] finetuned on the last layer (FT-Last) and on all layers (FT-All), Flan-T5-large [11] without fine-tuning, and CBM [18]. Notably, the first three are black-box models and CBM is interpretable by design.
As shown in Table 1, IP-CRR outperforms CXR-BERT (FT-Last), Flan-T5-large, and CBM in average precision (AP) and F1 score on the LO, CA, CM, and PN tasks, and achieves comparable performance on SD. However, compared to CXR-BERT (FT-All), IP-CRR matches its AP on LO, but underperforms on other tasks. We attribute this gap primarily to the CXR-BERT’s black-box nature, its more complex BERT-based architecture, and the fact that all its parameters are fine-tuned during training. In contrast, IP-CRR is implemented using a simple MLP with significantly fewer parameters, prioritizing interpretability.
Furthermore, in Figure 3, we evaluate AP as a function of the number of queries. For the LO task, IP-CRR achieves 0.95 AP with just 30 queries, and 0.97 with 100 – fewer than 20% of the total of 520 queries. Similar trends hold across other tasks, indicating that IP-CRR generally selects informative queries in its early iterations. Alternatively, we observe that IP-CRR converges to the maximum performance with only a small subset of queries. Compared to the baselines in Table 1, IP-CRR achieves better performance while using a smaller number of queries than CBMs, which use all available queries.
Next, in Figure 3, we plot the number of queries needed as the threshold on the posterior changes. The threshold on the posterior approximately measures the confidence the user requires the predictor to reach before terminating. We observe that the IP-CRR predictor requires fewer than 150 queries to achieve a confidence of 0.95 for most of tasks. This demonstrates that IP-CRR is able to find informative queries to make confident predictions.
@@ 4.3 Qualitative Analysis of Query-answer Chains
The core mechanics of using IP-CRR for interpretable predictions is the iterative process of selecting informative queries. This allows the user to observe how the posterior evolves as the number of queries increases. In Figure 4, we demonstrate two examples of query-answer chains obtained using IP-CRR. For the example on the left, our method first selects $q_{1}=$Opacification likely reflects atelectasis?”. From the report, the model infers that $q_{1}(x)$ is not present in the report. Then, it selects the query $q_{2}=$Patchy opacity at right lung base?”. Continuing in this fashion, after 9 queries, the model concludes that the patient has LO. The same sequential deduction can be observed in the example on the right, highlighting the interpretability IP-CRR offers when applied to CRRs.
Figure 4: Examples of query-answer chains from IP-CRR. Each row in the colored matrix represents the posterior $P(Y\mid q_{1:k}(x))$ after selecting the query $q_{k}$. Text in red means the answers is “no”, green means “yes”, and black means “unknown”. “N” or “P” means negative or positive prediction of the disease, respectively.
@@ 5 Conclusion
In this work, we have proposed an interpretable-by-design framework for CRR classification. We obtain a large set of queries by mining them from existing CRRs, and answer them using NLI. To evaluate the proposed framework, we conducted experiments on a large-scale CRR dataset and on multiple tasks, with both quantitative and qualitative results that illustrate interpretability and competitive performance. A current limitation is the reliance on a general-domain language model for query answering. We leave the use of stronger medical foundation models for generating more expressive query answers to future work.
{credits}
5.0.1 Acknowledgements
The authors thank Kaleab A. Kinfu, Uday Kiran Reddy Tadipatri and Prof. Dennis Parra for insightful discussions and feedback during the development of this work. Yuyan, Ryan, and René acknowledge the support from the Penn Engineering Dean’s Fellowship, Penn startup funds, and the NIH grant R01NS135613. Pablo was supported in Chile by ANID through iHEALTH (ICN2021_004), CENIA (FB210017), and the ANID Scholarship Program/Doctorado Becas Chile/2019-21191569.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.